Barlow Company manufactures three products-A, B, and C. The selling price, variable costs, and contribution margin for one unRequired 1 Required 2 Required 3 Required 4 ------------- Calculate the contribution margin per pound of the constraining resRequired 1 Required 2 Required 3 Required 4 Assuming that Barlow has unlimited demand for each of its three products, what isRequired 1 Required 2. Required 3 Required 4 Assuming that Barlows estimated customer demand is 600 units per product line,Required 1 Required 2 Required 3 Required 4 A foreign supplier could furnish Barlow with additional stocks of the raw materia
Answer #1
1 Particulars A B C
Direct material cost 18 90 27
Direct material pounds/unit 2 10 3
Contribution 36 90 36
Contribution margin per pound 18 9 12
2 Assuming, company has unlimited demand for all the products, Company will produce only product A since its contibution/pound is higher of the rest
Maximum production of product A = 6,300/2 = 2,100 units
Maximum contribution to company = 2,100 *36 = 75,600
3 Production
Particulars A B C
Contribution margin per pound 18 9 12
Rank I III II
Maximum demand                600                600                600
Raw material pounds required to produce to meet maximum demand            1,200            6,000            1,800
Allocation of raw material            1,200            3,300            1,800
Units to produce to get max profit                600                330                600
Contribution margin per unit                  36                  90                  36
Total contribution          21,600          29,700          21,600
4 Particulars A B C
Maximum demand                600                600                600
Units that can be produced with available raw materials                600                330                600
Balance                   -                  270                   -  
Raw materials required to produce additional units                   -              2,700                   -  
Extra contribution earned on producing additional units = 270* 90
                = 24,300
Maximum price that barlow could pay to foreign supplier = 24,300/2700
                    = 9/pound